# Emergent time-keeping mechanisms in a deep reinforcement learning agent performing an interval timing task

### v2

@@ Introduction
A recent research interest has emerged in drawing parallels between Deep Artificial Neural Networks (DNNs) and biological systems (Kanwisher et al.,, 2023; Richards et al.,, 2019). The similarities between the two are noteworthy because, DNNs exhibit emergent behavior and are not explicitly engineered to mimic biological systems. Unlike biological systems, which are a result of complex evolutionary adaptations, DNNs rely on relatively simple optimization techniques and architectures to solve problems. These two distinct systems converging to similar behaviors suggests important implications about the nature of the problem being solved. It may indicate that certain behaviors and representations are fundamentally useful to find a solution, and that problems presumed to require complex solutions might, in fact, be addressed through simpler means. Additionally, DNNs, with their flexibility for ablation studies and architectural manipulations, provide a powerful platform for systematically testing different hypotheses. Hence, identifying similarities between the two systems can facilitate better understanding of biological systems.
@@
Temporal processing is an extensively researched topic in both psychology and neuroscience (Wittmann,, 2009). However, the lack of a dedicated sensory organ for perceiving time, and scattered evidence regarding the involvement of multiple brain regions, makes it difficult to form a coherent understanding of temporal processing in humans and animals. Interval timing provides a framework to study these temporal processing mechanisms. The ability of humans and animals to successfully produce, reproduce, or categorize different temporal durations suggests the existence of an internal time-keeping mechanism. Recent research has begun exploring interval timing also in deep reinforcement learning (DRL) agents (Deverett et al.,, 2019; Lin et al.,, 2023). These studies additionally explore potential biological counterparts to the internal time-keeping mechanisms that emerge in the DRL agent performing interval timing tasks. They put forth a discussion about using advancements in Artificial Intelligence (AI) to enhance our understanding of temporal encodings in the brain.
@@
In this study, we investigate the internal time-keeping mechanisms of a DRL agent trained on a duration production task (Thönes and Stocker,, 2019). We show that the agent develops a robust internal time-keeping mechanism, characterized by oscillatory activation patterns, a motif commonly observed in biological systems. Furthermore, the agent’s timing strategy shares high-level functional similarities with the Striatal Beat Frequency (SBF) model, a biologically plausible model of interval timing (Matell and Meck,, 2000). Finally, the robust and internalized time-keeping mechanism developed in the agent is compared to certain aspects of the evolution of biological processes like circadian rhythms. Through this study, we aim to contribute to research that draws parallels between temporal processing in artificial agents and biological systems. It is our view that such efforts may provide alternative approaches to understanding temporal processing in humans and animals.
@@
The DRL agent architecture consisted of three primary components: a Convolutional Neural Network (CNN) to extract spatial features from each video frame, a Long Short-Term Memory (LSTM) network to process temporal dependencies across frames, and a linear layer to process the resulting representation and to compute the state value and next action (see Figure 1). The agent was implemented using the recurrent policy variant of the Proximal Policy Optimization (PPO) algorithm provided by the Stable-Baselines3 (SB3) library (Raffin et al.,, 2021). The reinforcement learning environment was a custom built environment developed using the Gymnasium framework (Towers et al.,, 2024). It was designed to deliver sequential video frames and handle action and rewards specific to the target interval time. The analysis conducted in this study focuses exclusively on the action network.
The agent was evaluated on a set of video sequences extracted from videos other than the training video. These test videos were categorized based on their Mean Absolute Frame Difference (MAFD), computed as described in Salih and George, (2020) and averaged over all frames for each video sequence (see Table 1). We use the average MAFD as a simplified metric to quantify the amount of visual information in a video sequence. A blank video, or one in which the same frame is repeated throughout (i.e., constant frame video), yields a MAFD of zero and thus contains minimal visual information. In contrast, videos with changing frames exhibit higher MAFD values and thus contain more visual information. It is worth noting that, a video sequence composed of random frames exhibits even higher MAFD values, despite lacking meaningful visual information. Thus, although high MAFD reflects greater pixel-level changes between video frames, it does not necessarily correspond to meaningful visual content. However, since the agent’s primary objective is not to process visual information as such, but to produce specific intervals, we assume that higher MAFD corresponds to greater visual information from the agent’s perspective, with the caveat that this metric captures only low-level variability.
@@
The agent successfully managed to learn the timing task and was able to produce the target interval consecutively throughout the video sequence. Analysis of the agent’s internal states revealed that the LSTM layer played a crucial role in the development of an intrinsic timekeeping mechanism. This was expected since the LSTM network, which is a type of artificial Recurrent Neural Network (RNN), is designed to extract temporal features from sequential data (Hochreiter and Schmidhuber,, 1997; Yu et al.,, 2019). An analysis of the LSTM layer’s hidden state activations reveals emergent high amplitude oscillatory patterns with frequencies correspond to the target interval.
@@
Interestingly, a prior study involving a DRL agent trained for time reproduction (with spatial input) did not report such oscillatory patterns in the PCA of the LSTM cell states (Deverett et al.,, 2019). The authors concluded that the cell state units in the LSTM exhibit counter-like behavior. However, in Lin et al.,, 2023, for certain conditions, oscillating activity has been reported in the RNN units of a DRL agent (with non-spatial inputs) for a duration comparison and an episodic working memory task. Additionally, oscillatory patterns have been observed in a continuous time recurrent neural networks (CTRNNs) optimized using genetic algorithm to perform multiple timing tasks Maniadakis and Trahanias, (2015). However, unlike the LSTM or RNN layer of the DRL agents, CTRNNs are known for their biologically plausible interpretation Yu et al., (2017), Beer, (1995). Thus, findings from current and prior studies suggests that oscillatory dynamics may serve as a fundamental mechanism of time-keeping for specific types of interval timing tasks in artificial agents, regardless of the biological plausibility of their architectures.
@@
Emergence of oscillatory patterns in the DRL agent’s network is noteworthy because oscillations are a known characteristic of various types of biological neurons (Heltberg et al.,, 2021; Buzsáki et al.,, 2013), including those involved in temporal processing. Importantly, the DRL agent architecture was not designed to be biologically plausible and it was not provided with any external timing cues or explicitly encoded neural dynamics. Despite this, it successfully learned to perform the timing task, and, remarkably, the hidden state activations of the network converged to an oscillatory pattern with frequencies corresponding to the target interval. These findings suggest that, despite its non-biological architecture, the DRL agent converges to use biologically plausible mechanisms, specifically, oscillatory dynamics, for time-keeping in a dedicated time production task.
@@
A biologically plausible model of interval timing called the Striatal Beat Frequency model (SBF), describes interval timing in terms of coincidence detection in neural oscillations Matell and Meck, (2000). According to this model, the striatal spiny neurons function as coincidence detectors of cortical neural oscillations. These coincidence detectors fire when a set amount of coincidental input activity (i.e., coincident oscillations of the cortical neurons) is reached. According to Buhusi and Meck, (2005) and Matell and Meck, (2000), for tasks involving the production of a learned interval, there is a burst of dopaminergic activity at the trial onset. This triggers synchronization in oscillation of the cortical neurons. Additionally, another burst in activity of dopaminergic neurons occurs at the expected time of reward which corresponds to the target interval. This causes an update in the weights of corticostriatal synapses and consequently triggers the firing of the striatal spiny neurons (or the coincidental detectors). As a result, the firing rate of these coincidence detectors can encode information about a previously learned interval. Two high-level functional abstractions of the SBF model are considered in this study to explore whether the DRL agent’s emergent time-keeping mechanism resembles the biologically plausible timing model.
@@
A potentially related phenomena can be observed in the evolution of circadian rhythms in biological systems. Circadian clocks in organisms are believed to have developed as a strategy of photosynthesizing organisms to cope with day and night fluctuations from the environment Hut and Beersma, (2011). These organisms developed oscillating molecular behaviors (typically with periods close to 24 hours) in order to develop different molecular processes corresponding to daylight and darkness. Notably, in some species, these circadian rhythms persist even in the absence of external cues, such as in constant darkness Seki and Ito, (2022). This highlights how internal mechanisms, initially shaped by environmental interactions in biological systems, can become self-sustaining. We hypothesize that, the DRL agent’s behavior related to first learning a task by interacting with its environment, and later internalizing it to be able to perform the task irrespective of the environments input has some parallels with the evolution of biological processes like circadian clocks.
@@
Table 1: Reward across videos with different levels of Mean Absolute Frame Difference (MAFD) (Salih and George,, 2020) averaged across all frames of the video sequence. The maximum achievable reward per episode is 5. While the agent generally performs well, its performance is slightly affected for video sequences with higher MAFD. Average frame difference of zero corresponds to a blank or constant frame video sequence and the average frame difference of 34.7 corresponds to a random video sequence.
@@ Discussion
The findings from this study, in conjunction with those of Deverett et al., (2019); Lin et al., (2023), suggest several implications for interval timing tasks. Specifically, the temporal representation required for successfully completing these tasks appears to be extractable by an artificial neural network from its input. This could imply the involvement of lower-level cognitive processes (König et al.,, 2013) in the execution of these tasks. This observation further supports the distinction made in time perception between duration judgment (i.e., the ability to estimate or reproduce time intervals) and the passage of time (i.e., how time feels) (Thönes and Stocker,, 2019). Specifically, while an artificial agent can learn to perform interval timing tasks with high accuracy, evaluating the subjective passage of time (whether time feels fast or slow), remains a significantly more challenging problem for such systems. This disparity may indicate the differing levels of cognitive complexity involved in the two tasks.
This study has several limitations. The parallels drawn between the agent’s behavior and biological systems, namely the resemblance to the SBF model of timing and the evolution of circadian rhythms, are preliminary and require further investigation. Given the complexity of biological systems, such analogies may overlook critical nuances. Two distinct systems arriving at similar solutions may not always imply shared mechanisms. This is because, the same problem can often be solved through multiple, distinct optimization pathways. Additionally, our analysis focused solely on the action network, whereas previous work has examined the value network and reported ramping-cell-like activity for temporal tasks (Lin et al.,, 2023). Future work could extend this study by analyzing the value network in our agent to uncover additional biologically plausible temporal patterns in the artificial neurons. Finally, this work represents an initial step towards using artificial agents to study temporal processing in biological systems. Understanding the implications of these similarities and translating them into meaningful biological insights remains a long-term challenge.

### v3

@@ Introduction
A recent research interest has emerged in drawing parallels between Deep Artificial Neural Networks (DNNs) and biological systems (Kanwisher et al., 2023; Richards et al., 2019). The similarities between the two are noteworthy because, DNNs exhibit emergent behavior and are not explicitly engineered to mimic biological systems. Unlike biological systems, which are a result of complex evolutionary adaptations, DNNs rely on relatively simple optimization techniques and architectures to solve problems. These two distinct systems converging to similar behaviors suggests important implications about the nature of the problem being solved. It may indicate that certain behaviors and representations are fundamentally useful to find a solution, and that problems presumed to require complex solutions might, in fact, be addressed through simpler means. Additionally, DNNs, with their flexibility for ablation studies and architectural manipulations, provide a powerful platform for systematically testing different hypotheses. Hence, identifying similarities between the two systems can facilitate better understanding of biological systems.
@@
Temporal processing is an extensively researched topic in both psychology and neuroscience (Wittmann, 2009). However, the lack of a dedicated sensory organ for perceiving time, and scattered evidence regarding the involvement of multiple brain regions, makes it difficult to form a coherent understanding of temporal processing in humans and animals. Interval timing provides a framework to study these temporal processing mechanisms. The ability of humans and animals to successfully produce, reproduce, or categorize different temporal durations suggests the existence of an internal time-keeping mechanism. Recent research has begun exploring interval timing also in deep reinforcement learning (DRL) agents (Deverett et al., 2019; Lin et al., 2023). These studies additionally explore potential biological counterparts to the internal time-keeping mechanisms that emerge in the DRL agent performing interval timing tasks. They put forth a discussion about using advancements in Artificial Intelligence (AI) to enhance our understanding of temporal encodings in the brain.
@@
In this study, we investigate the internal time-keeping mechanisms of a DRL agent trained on a duration production task (Thönes and Stocker, 2019). We show that the agent develops a robust internal time-keeping mechanism, characterized by oscillatory activation patterns, a motif commonly observed in biological systems. Furthermore, the agent’s timing strategy shares high-level functional similarities with the Striatal Beat Frequency (SBF) model, a biologically plausible model of interval timing (Matell and Meck, 2000). Finally, the robust and internalized time-keeping mechanism developed in the agent is compared to certain aspects of the evolution of biological processes like circadian rhythms. Through this study, we aim to contribute to research that draws parallels between temporal processing in artificial agents and biological systems. It is our view that such efforts may provide alternative approaches to understanding temporal processing in humans and animals.
@@
The DRL agent architecture consisted of three primary components: a Convolutional Neural Network (CNN) to extract spatial features from each video frame, a Long Short-Term Memory (LSTM) network to process temporal dependencies across frames, and a linear layer to process the resulting representation and to compute the state value and next action (see Figure 1). The agent was implemented using the recurrent policy variant of the Proximal Policy Optimization (PPO) algorithm provided by the Stable-Baselines3 (SB3) library (Raffin et al., 2021). The reinforcement learning environment was a custom built environment developed using the Gymnasium framework (Towers et al., 2024). It was designed to deliver sequential video frames and handle action and rewards specific to the target interval time. The analysis conducted in this study focuses exclusively on the action network.
The agent was evaluated on a set of video sequences extracted from videos other than the training video. These test videos were categorized based on their Mean Absolute Frame Difference (MAFD), computed as described in Salih and George (2020) and averaged over all frames for each video sequence (see Table 1). We use the average MAFD as a simplified metric to quantify the amount of visual information in a video sequence. A blank video, or one in which the same frame is repeated throughout (i.e., constant frame video), yields a MAFD of zero and thus contains minimal visual information. In contrast, videos with changing frames exhibit higher MAFD values and thus contain more visual information. It is worth noting that, a video sequence composed of random frames exhibits even higher MAFD values, despite lacking meaningful visual information. Thus, although high MAFD reflects greater pixel-level changes between video frames, it does not necessarily correspond to meaningful visual content. However, since the agent’s primary objective is not to process visual information as such, but to produce specific intervals, we assume that higher MAFD corresponds to greater visual information from the agent’s perspective, with the caveat that this metric captures only low-level variability.
@@
The agent successfully managed to learn the timing task and was able to produce the target interval consecutively throughout the video sequence. Analysis of the agent’s internal states revealed that the LSTM layer played a crucial role in the development of an intrinsic timekeeping mechanism. This was expected since the LSTM network, which is a type of artificial Recurrent Neural Network (RNN), is designed to extract temporal features from sequential data (Hochreiter and Schmidhuber, 1997; Yu et al., 2019). An analysis of the LSTM layer’s hidden state activations reveals emergent high amplitude oscillatory patterns with frequencies correspond to the target interval.
@@
Interestingly, a prior study involving a DRL agent trained for time reproduction (with spatial input) did not report such oscillatory patterns in the PCA of the LSTM cell states (Deverett et al., 2019). The authors concluded that the cell state units in the LSTM exhibit counter-like behavior. However, in Lin et al., 2023, for certain conditions, oscillating activity has been reported in the RNN units of a DRL agent (with non-spatial inputs) for a duration comparison and an episodic working memory task. Additionally, oscillatory patterns have been observed in a continuous time recurrent neural networks (CTRNNs) optimized using genetic algorithm to perform multiple timing tasks Maniadakis and Trahanias (2015). However, unlike the LSTM or RNN layer of the DRL agents, CTRNNs are known for their biologically plausible interpretation Yu et al. (2017), Beer (1995). Thus, findings from current and prior studies suggests that oscillatory dynamics may serve as a fundamental mechanism of time-keeping for specific types of interval timing tasks in artificial agents, regardless of the biological plausibility of their architectures.
@@
Emergence of oscillatory patterns in the DRL agent’s network is noteworthy because oscillations are a known characteristic of various types of biological neurons (Heltberg et al., 2021; Buzsáki et al., 2013), including those involved in temporal processing. Importantly, the DRL agent architecture was not designed to be biologically plausible and it was not provided with any external timing cues or explicitly encoded neural dynamics. Despite this, it successfully learned to perform the timing task, and, remarkably, the hidden state activations of the network converged to an oscillatory pattern with frequencies corresponding to the target interval. These findings suggest that, despite its non-biological architecture, the DRL agent converges to use biologically plausible mechanisms, specifically, oscillatory dynamics, for time-keeping in a dedicated time production task.
@@
A biologically plausible model of interval timing called the Striatal Beat Frequency model (SBF), describes interval timing in terms of coincidence detection in neural oscillations Matell and Meck (2000). According to this model, the striatal spiny neurons function as coincidence detectors of cortical neural oscillations. These coincidence detectors fire when a set amount of coincidental input activity (i.e., coincident oscillations of the cortical neurons) is reached. According to Buhusi and Meck (2005) and Matell and Meck (2000), for tasks involving the production of a learned interval, there is a burst of dopaminergic activity at the trial onset. This triggers synchronization in oscillation of the cortical neurons. Additionally, another burst in activity of dopaminergic neurons occurs at the expected time of reward which corresponds to the target interval. This causes an update in the weights of corticostriatal synapses and consequently triggers the firing of the striatal spiny neurons (or the coincidental detectors). As a result, the firing rate of these coincidence detectors can encode information about a previously learned interval. Two high-level functional abstractions of the SBF model are considered in this study to explore whether the DRL agent’s emergent time-keeping mechanism resembles the biologically plausible timing model.
@@
A potentially related phenomena can be observed in the evolution of circadian rhythms in biological systems. Circadian clocks in organisms are believed to have developed as a strategy of photosynthesizing organisms to cope with day and night fluctuations from the environment Hut and Beersma (2011). These organisms developed oscillating molecular behaviors (typically with periods close to 24 hours) in order to develop different molecular processes corresponding to daylight and darkness. Notably, in some species, these circadian rhythms persist even in the absence of external cues, such as in constant darkness Seki and Ito (2022). This highlights how internal mechanisms, initially shaped by environmental interactions in biological systems, can become self-sustaining. We hypothesize that, the DRL agent’s behavior related to first learning a task by interacting with its environment, and later internalizing it to be able to perform the task irrespective of the environments input has some parallels with the evolution of biological processes like circadian clocks.
@@
Table 1: Reward across videos with different levels of Mean Absolute Frame Difference (MAFD) (Salih and George, 2020) averaged across all frames of the video sequence. The maximum achievable reward per episode is 5. While the agent generally performs well, its performance is slightly affected for video sequences with higher MAFD. Average frame difference of zero corresponds to a blank or constant frame video sequence and the average frame difference of 34.7 corresponds to a random video sequence.
@@ Discussion
The findings from this study, in conjunction with those of Deverett et al. (2019); Lin et al. (2023), suggest several implications for interval timing tasks. Specifically, the temporal representation required for successfully completing these tasks appears to be extractable by an artificial neural network from its input. This could imply the involvement of lower-level cognitive processes (König et al., 2013) in the execution of these tasks. This observation further supports the distinction made in time perception between duration judgment (i.e., the ability to estimate or reproduce time intervals) and the passage of time (i.e., how time feels) (Thönes and Stocker, 2019). Specifically, while an artificial agent can learn to perform interval timing tasks with high accuracy, evaluating the subjective passage of time (whether time feels fast or slow), remains a significantly more challenging problem for such systems. This disparity may indicate the differing levels of cognitive complexity involved in the two tasks.
This study has several limitations. The parallels drawn between the agent’s behavior and biological systems, namely the resemblance to the SBF model of timing and the evolution of circadian rhythms, are preliminary and require further investigation. Given the complexity of biological systems, such analogies may overlook critical nuances. Two distinct systems arriving at similar solutions may not always imply shared mechanisms. This is because, the same problem can often be solved through multiple, distinct optimization pathways. Additionally, our analysis focused solely on the action network, whereas previous work has examined the value network and reported ramping-cell-like activity for temporal tasks (Lin et al., 2023). Future work could extend this study by analyzing the value network in our agent to uncover additional biologically plausible temporal patterns in the artificial neurons. Finally, this work represents an initial step towards using artificial agents to study temporal processing in biological systems. Understanding the implications of these similarities and translating them into meaningful biological insights remains a long-term challenge.